OpenStudy (anonymous):

(-2,-21)(3,4)(-7,-46), they satisfy a linear relationship, what is the value of x when y= 14?

OpenStudy (anonymous):

I got an answer earlier, but I didn't understand how you got from a to b :O

OpenStudy (anonymous):

Pick any two of these points. They must be located on some line, which can be described using an equation of form: y = ax + b Using x and y values from the points you selected and a system of 2 equations like the one above, find a and b that satisfy it. Then solve using y = 14.

OpenStudy (anonymous):

I'm still a little lost. I understand that you do that, however when you get: -21=-2a+b and 4=3a+b how do you then get -25= -5+b?

OpenStudy (anonymous):

A small mistake there: -21 = -2a + b

OpenStudy (anonymous):

Oh yes. Can you explain how you get there? Am I supposed to automatically assume that you subtract in this eqution?

OpenStudy (anonymous):

Sure. Let's start with the points you picked. We get: -21 = -2a + b 4 = 3a + b Let's calculate 'b' using the first equation: b = -21 + 2a Now we put it in the 2nd one: 4 = 3a + (-21 + 2a) Simplification gives us: 4 = 3a - 21 + 2a 5a = 25 a = 5 Now we use the calculated value of 'a' in the first equation: b = -21 + 2a b = -21 + 2 * 5 b = - 21 + 10 b = -10 So the line you're looking for is given by the equation: y = 5x - 10

OpenStudy (anonymous):

So you're subsituting in the equations to find the ultimate answer?

OpenStudy (anonymous):

I used the substitution method, yes.

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

Welcome!